OpenStudy (anonymous):

The area of a rectangle can be found by multiplying the length by the width. A rectangular postage stamp has a length of nine-tenths inch and a width of five-sixths inch. What is the area of the stamp, in square inches? three-fourths square inches one-seventeenth square inches one-half square inches one and one-fourth square inches

OpenStudy (anonymous):

10 and 6 have a common multipliple of 30 so..\[\frac{ 27 }{ 30 }*\frac{ 25 }{ 30 }=\frac{ 675 }{ 900 }=.75=\frac{ 3 }{ 4 }\]

OpenStudy (anonymous):

\[or \frac{ 9 }{ 10 }*\frac{ 5 }{ 6 }=\frac{ 45 }{ 60 }=\frac{ 3 }{ 4 }\]